Sunflowers exude a type of sap on a regular schedule each day, suggesting the operation of a biological clock. The following parameters were measured.

TAU OF EXUDATION CYCLE (HRS)VOLUME OF EXUDATION IN 24 HRS GROWTH TEMPERATURE (°C) 20°C plants 30°C plants (MICRO-LITERS/PLANT) 89 25.3 223 24.1All plants were grown in 12L:12D before the experiment, and they were kept under constant dim light during data collection.

A) Calculate the Q10 for each of the two different variables. For the period data, calculate Q10 as a simple ratio with the data from the lower temperature as the numerator and the data from the higher temperature as the denominator. For the volume data, calculate Q10 as a simple ratio with the data from the higher temperature as the numerator and the data from the lower temperature as the denominator. For each variable, write two sentences explaining whether or not you find evidence of temperature dependence and of temperature compensation.

B) Explain why the numerators and denominators switch places in the two different types of sunflower measurements ... you can figure this out if you think of the mathematical relationship between period and tau.
Answer #1

A) Q 10 of TAU = 25.3/24.1 = 1.04

Q10 of volume = 223/89 = 2.5

For both the values there is a significant difference in the values upon increase in temperature by 10 degree Celsius.

B) Q10 values denotes increasing thermal difference the circadian rhythm is inversely proportional to the protein expressed.